OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Passing through (-5, -7) and (-8, -6)

OpenStudy (danjs):

calculate the slope by using those two points first...

OpenStudy (danjs):

\[(y-y _{1}) = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }*(x - x_{1})\]

OpenStudy (danjs):

Use those two points, (x1 , y1) and (x2,y2), plug in there

OpenStudy (anonymous):

Slope = -0.3 θ = 161 ΔX = -3 ΔY = 1

OpenStudy (astrophysics):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]